OpenStudy (anonymous):

slope=1/2 through (-1,-3)

OpenStudy (rulnick):

y = (1/2) x - 5/2

OpenStudy (rulnick):

notice that (-1,-3) works: -3 = (1/2) (-1) - 5/2

OpenStudy (anonymous):

could you give me an example how you got the answer

OpenStudy (rulnick):

set y = (1/2) x + b because you know the slope (1/2) but not the intercept (b) then plug in x and y (x=-1, y=-3) solve for b done.